# On the Estimation of bivariate Conditional Transition Rates

### v1

@@ 6 Uniform Convergence in Actuarial Estimation
The proof of this theorem uses the same ideas already used for the continuity of the two-dimensional cash-flow. We use the fact that $A_{i}$ and $a_{ij}$ both have finite variation and the fact, that integrals of these forms are Hadamard differentiable, see Lemma A.3. ∎
@@ 7 Concluding Remarks
The use of analytical techniques well-known from bivariate Kaplan-Meier estimation and conditional Aalen-Johansen estimation leads to uniform convergence of both estimators. We obtain a blueprint for the estimation of bivariate conditional transition rates and transition probabilities from right-censored data.
Future research endeavours could address the question of asymptotic normality and bootstrapping. Additionally, an extension of the uniform convergence to the entire plane, possibly via martingale techniques, is of interest.
Finally, this study closes a gap between techniques in survival analysis and premium and reserve calculation for insurance contracts. In the case of an insurance contract with scaled payments, we can construct estimators for the first moment of all future payments. This is the basis of any type of risk assessment, where the state process of the insured follows the non-restrictive non-Markov model. Furthermore, estimating bivariate conditional transition rates and transition probabilities allows the estimation of second moments of future liabilities with a canonical one-dimensional cash-flow in the non-Markov model. Altogether, this can be used to evaluate the widespread use of Markov assumptions in the reserving and calculation for insurance portfolios.
@@ Appendix A Appendix
For a monotonously increasing real-matrix valued function $\Lambda^{(\epsilon)}$ and a function $\phi^{(\epsilon)}(\boldsymbol{t})$, the equation
$Y(\boldsymbol{t})=\phi^{(\epsilon)}(\boldsymbol{t})+\int\displaylimits_{(s,\boldsymbol{t}]}Y(\boldsymbol{u}^{-})\Lambda^{(\epsilon)}(\mathrm{d}\boldsymbol{u}),$ (A.1)
@@
$\begin{split}\tilde{P}^{(\epsilon)}(\boldsymbol{t})&=\phi^{(\epsilon)}(\boldsymbol{t})+\int\displaylimits_{(s,\boldsymbol{t}]}\phi^{(\epsilon)}(\boldsymbol{u}^{-})\Lambda^{(\epsilon)}(\mathrm{d}\boldsymbol{u})\mathcal{P}((\boldsymbol{u},\boldsymbol{t}],\Lambda^{(\epsilon)}),\end{split}$ (A.2)
@@
For monotonously increasing functions $A,B:\mathbb{R}^{2}\rightarrow\mathbb{R}^{n\times n}$ we have the following equation
@@
Additionally, we need a two-dimensional generalisation of the well known Campbell theorem, see for example Milbrodt & Helbig [19]. We add the proof for this generalisation because in standard literature only the one-dimensional case is used.
@@
For the estimation of integrals, we use the so called one and two-dimensional variation of a process. We will do a short introduction here.

### v2

@@ 6 Uniform Convergence in Actuarial Estimation
The proof of this theorem uses the same ideas already used for the continuity of the two-dimensional cash flow. We use the fact that $A_{i}$ and $a_{ij}$ both have finite variation and the fact, that integrals of these forms are Hadamard differentiable, see Lemma A.3. ∎
7 Numerical Example
We consider a numerical example, where the bivariate conditional transition rates and transition probabilities arise naturally in the conditional expected values of a future cash flow. The focus of the numerical study is to illustrate how our two-dimensional cash flows from Section 6 can be applied in practice and how censoring affects the estimations. This is not intended to be a full statistical analysis, but merely an exploratory example.
The basic setup is similar to that used for the numerical study in the paper by Buchardt and Møller [9]. As a technical basis, we consider a survival model, i.e. a Markov model with the two states alive (1) and dead (4). The modelled insurance contract starts at the age of 40 and the retirement age is 65. The insured pays a lump sum of 100,000 at the beginning of the contract and an annual premium of 10,000 until the insured reaches the age of 65 and receives annual pension payments. All payments cease when the insured dies or reaches the age of 100. We get the following sojourn payment function:
$B_{1}(\operatorname{dt})$ $:=(-10,000\mathds{1}_{\{t<65\}}+p\mathds{1}_{\{100\geq t\geq 65\}})\operatorname{dt},$
where $p$ is the annuity calculated on the technical basis according to the equivalence principle. The technical basis is generally used to price insurance contracts, while the extended model of the market basis is used for the insurers’ balance sheets. The technical basis uses the transition intensity
$\mu^{\star}_{14}(t):=0.005+10^{(5.728-10+0.038t)},$
and an interest rate of $0\%$. The market basis consists of four states, with states (1) and (4) denoting alive and dead states, mirroring the technical basis. In addition, state (2) denotes the free policy option and state (3) denotes surrender, where free policy means that premiums are waived and benefits are reduced accordingly. Surrender means a complete cessation of both benefits and payments with a final transitional payment, usually in the form of the current reserve. We therefore introduce two options that model policyholder behaviour for the market basis. The additional payment functions are modelled as follows:
$b_{13}(t)$ $=V_{1}(t),$
$b_{23}(t,u)$ $=V_{1}^{+}(t)\rho(u),$
for the transition payments, where $V_{1}^{+}(t)$ is the technical reserve of the pension payments, and
$B_{2}(\operatorname{dt},u)$ $:=\rho(u)p\mathds{1}_{\{100\geq t\geq 65\}}\operatorname{dt},$
for the sojourn payments, where $p$ is the premium that is paid in state (1). The factor $\rho(u)$ defines the reduction of the benefits in the free policy state (2). It is calculated by
$\rho(u)=\frac{A_{1}^{+}(u)-A_{1}^{-}(u)}{A_{1}^{+}(u)},$
where $A_{1}^{-}(t)$ is the technical reserve of premium payments. The dependence of the payments functions on $u$ models the dependence on the transition time to state (2). For this reason, bivariate transition rates are natural in this application. See the work of Ahmad et al. in [3] for the theoretical background of free policy option calculations.
We model the market basis as a semi-Markov model with transition intensities
$\mu_{14}(t)$ $:=\mu^{\star}_{14}(t),$
$\mu_{24}(t)$ $:=\mu^{\star}_{14}(t),$
$\mu_{12}(t)$ $:=0.1\mathds{1}_{\{t\leq 65\}}$
$\mu_{13}(t)$ $:=0.05\mathds{1}_{\{t\leq 65\}}$
$\mu_{23}(t,u)$ $:=(0.05+0.2\mathds{1}_{\{u\in[\frac{1}{2},\frac{5}{2})\}})\mathds{1}_{\{t\leq 65\}},$
and an interest rate of $3\%$ annually. We estimate the conditional expected future cash flow of the market basis in an as-if Markov model, i.e. the landmark is $\xi=Z(40)$ and since 40 is the starting point of the insurance contract, we assume that each insured is in state (1) at $s=40$. For an as-if Markov model and a cash flow with scaled payments such as the free policy option, Bathke and Christiansen in [5] found a two-dimensional canonical representation for the resulting conditional expected future cash flow. Adapting it to this example, we get
$\begin{split}\boldsymbol{A}_{1}(t)&=\int\displaylimits_{(40,t]}\frac{\kappa(40)}{\kappa(u)}p_{1,1}(u^{-})B_{1}(\mathrm{d}u)+\int\displaylimits_{(40,t]}\frac{\kappa(40)}{\kappa(u)}b_{13}(u)p_{1,1}(u^{-})\Lambda_{1,(1,3)}(\mathrm{d}u)\\ &\quad+\int\displaylimits_{(40,t]}\;\int\displaylimits_{(40,65]\times(40,t)}\frac{\kappa(40)}{\kappa(t)}\rho(u_{1})\bigg{(}\boldsymbol{p}_{1,(1,1)}(\boldsymbol{u^{-}})\boldsymbol{\Lambda}_{1,(1,1)(2,2)}(\mathrm{d}\boldsymbol{u})\\ &\hskip 142.26378pt-\sum_{k\in\{3,4\}}\boldsymbol{p}_{1,(1,2)}(\boldsymbol{u}^{-})\boldsymbol{\Lambda}_{1,(1,2)(2,k)}(\mathrm{d}\boldsymbol{u})\bigg{)}p\mathds{1}_{\{t\geq 65\}}\mathrm{d}t\\ &\quad+\int\displaylimits_{(40,t]\times(40,t]}\frac{\kappa(s)}{\kappa(u_{1})}\rho(u_{1})b_{23}(u_{2},u_{1})\boldsymbol{p}_{1,(1,2)}(\boldsymbol{u^{-}})\boldsymbol{\Lambda}_{1,(1,2)(2,3)}(\mathrm{d}\boldsymbol{u}),\end{split}$ (7.1)
where $t\leq 100$ and $\kappa(\cdot)$ is the value of a deterministic cash account. The conditional expected value of this cash flow is an example where the estimation using the as-if Markov approach is natural, because we have jumps from state (2) to state (3) or (4) in the cash flow that are dependent on the jump from state (1) to state (2). Such dependencies are the strength of the conditional bivariate estimation.
We introduce some technical background to the estimation of this expected future cash flow in the as-if Markov model. The estimation of the conditional bivariate transition rates and transition probabilities for the market basis can be simplified. We can reduce the required transition rates to $\Lambda_{(1,2),(2,4)}$,$\Lambda_{(1,2),(2,3)}$, $\Lambda_{(1,2),(1,2)}$, $\Lambda_{(1,4),(1,4)}$ and $\Lambda_{(1,3),(1,3)}$ where the last two are only needed to estimate the transition probability $p_{11}$. This reduction originates from the reduced number of possible jumps in the market basis. The required transition rates can also be reduced. We only need $p_{11},p_{12}$ and $p_{13}$. This greatly reduces computation time and space.
Counting the number of steps required to compute the transition rates and probabilities, we get a time complexity of order $\mathcal{O}(n^{2})$ and a space complexity of order $\mathcal{O}(n^{2})$. Intuitively, we get this complexity because there are $\mathcal{O}(n^{2})$ time points on the time grid given by the data, and we need to recursively compute the estimator on this grid.
To compute the plug-in estimator, we can use a type of differential equation, in this case a simple recursive formula due to the finiteness of the grid points. This reduces the computational complexity for the computation of the conditional expected future cash flow from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n^{2})$. We use the decomposition
$\boldsymbol{A}_{1}({P}_{1}^{(n,\epsilon)},\boldsymbol{P}_{1}^{(n,\epsilon)},\Lambda_{1}^{(n,\epsilon)},\boldsymbol{\Lambda}_{1}^{(n,\epsilon)})(t)-\boldsymbol{A}_{1}(0)=:C_{1}(t)+C_{2}(t)+C_{3}(t),\quad t\geq 40,$
where $C_{1}$ contains the first two, $C_{2}$ the third and $C_{3}$ the fourth summand of the equation (7.1), with the estimated transition rates and transition probabilities plugged in. Now all the three summands can be calculated recursively. Let $\{t_{n}\}_{n\in I}$ be the the set of all jump points defining the time grid imposed by the data. We start with $t_{0}=40$ and $C_{1}(40)=C_{2}(40)=C_{3}(40)=0$. We get
$C_{1}(t_{n})=C_{1}(t_{n-1})+\int_{t_{n-1}}^{t_{n}}p^{(n,\epsilon)}_{1,j}(s-)\Big{(}B_{1}(\mathrm{d}s)+b_{13}(s)\Lambda^{(n,\epsilon)}_{1,13}(\mathrm{d}s)\Big{)}.$
For the second one, we get
$C_{2}(t_{n})$ $=C_{2}(t_{n-1})+\int_{(40,t_{n-1}-)}^{(65,t_{n}-)}\int_{u_{2}}^{t_{n}}B_{2}(\mathrm{d}s,u_{1})\bigg{(}\boldsymbol{P}^{(n,\epsilon)}_{1,(1,1)}(\boldsymbol{u-})\boldsymbol{\Lambda}^{(n,\epsilon)}_{1,(1,1)(2,2)}(\mathrm{d}\boldsymbol{u})$
$\hskip 176.40707pt-\sum_{k\in\{3,4\}}\boldsymbol{P}^{(n,\epsilon)}_{1,(1,2)}(\boldsymbol{u}-)\boldsymbol{\Lambda}^{(n,\epsilon)}_{1,(1,2)(2,k)}(\mathrm{d}\boldsymbol{u})\bigg{)}$
$\hskip 52.63759pt+\int_{t_{n-1}}^{t_{n}}p\mathds{1}_{\{t\geq 65\}}\mathrm{d}(s)\int_{(40,40)}^{(65,t_{n-1}-)}\rho(u_{1})\bigg{(}\boldsymbol{P}^{(n,\epsilon)}_{1,(1,1)}(\boldsymbol{u-})\boldsymbol{\Lambda}^{(n,\epsilon)}_{1,(1,1)(2,2)}(\mathrm{d}\boldsymbol{u})$
$\hskip 176.40707pt-\sum_{k\in\{3,4\}}\boldsymbol{P}^{(n,\epsilon)}_{1,\boldsymbol{j}}(\boldsymbol{u}-)\boldsymbol{\Lambda}^{(n,\epsilon)}_{1,(1,2)(2,k)}(\mathrm{d}\boldsymbol{u})\bigg{)}.$
The last integral in the formula for $C_{2}$ can again be calculated recursively in every step. For $C_{3}$, we get
$C_{3}(t_{n})=C_{3}(t_{n-1})$ $+\int_{(40,t_{n-1})}^{(65,t_{n})}b_{23}(u_{2},u_{1})\boldsymbol{P}^{(n,\epsilon)}_{1,(12)}(\boldsymbol{u-})\boldsymbol{\Lambda}^{(n,\epsilon)}_{1,(1,2)(2,3)}(\mathrm{d}\boldsymbol{u}).$
We estimate in R [27] by generating artificial data following the semi-Markov model described as the market basis using the package AalenJohansen by Bladt and Furrer [6].
We simulate $5,000$ independent and identically distributed realisations for the defined insurance contract with scaled payments. We use $\operatorname{Unif}(65,120)$ distributed random variables as right censoring.
(a) $n=1000$
(b) $n=5000$
Figure 7.1: Expected future cash flow $\boldsymbol{A}_{1}(t)$ with time horizon $T=100$ in the as-if Markov model estimated with $n=1000$ and $n=5000$ censored individuals, the plug-in estimator and the landmark Nelson-Aalen and Aalen-Johansen estimators versus Monte Carlo estimation with $n=10000$ uncensored individuals.
To illustrate the numerical feasibility and strength of the plug-in estimator, we plot in figures 7.1a and 7.1b the estimate for the conditional expected future cash flow of the insurance contract in an as-if Markov model on the market basis for $n=1000$ and $n=5000$ individuals, together with a Monte Carlo simulation of the same conditional expected future cash flow without censoring and $10,000$ individuals for comparison. We see that the bivariate estimator, with $n=5000$, estimates the conditional expected future cash flow of the particular insurance product similarly to the Monte Carlo estimator, which is using uncensored data and more individuals. Thus, the estimator appears to be a good fit for this type of estimation problem.
A challenge for the interpretation of the estimator $P^{(n,\epsilon)}_{z,(jk)}$ is that in the censored case the bivariate estimator may not satisfy the expected monotonicity and positivity properties. This property is very similar to the findings in [13] where they found the Kaplan-Meier estimator on the plane to be non-monotonic. This explains the higher volatility observed in the plug-in estimator for $n=1000$ but does not change the fact that the estimator consistently estimates the conditional expected future cash flow, even when censoring is involved. Moreover, the estimators are easy to apply, since we do not need to solve any continuous differential equations, but were able to obtain recursive formulas for the estimators at each level.
We conclude that, despite numerical challenges, the bivariate estimator can be successfully applied to the estimation of conditional expected values of future cash flows in non-Markov models. The bivariate landmark Nelson-Aalen and Aalen-Johansen estimators, combined with the plug-in estimators for conditional expected future cash flows, provide a versatile tool because the theory applies not only to insurance contracts with contract modifications, but also to other cash flows that have a two-dimensional representation.
@@ 8 Concluding Remarks
The use of analytical techniques well-known from bivariate Kaplan-Meier estimation and conditional Aalen-Johansen estimation leads to uniform convergence of both estimators. We obtain a blueprint for the estimation of bivariate conditional transition rates and transition probabilities from right censored data.
Future research endeavours could further address the question of asymptotic normality and bootstrapping. Additionally, an extension of the uniform convergence to the entire plane, possibly via martingale techniques, is of interest.
Finally, this study closes a gap between techniques in survival analysis and premium and reserve calculation for insurance contracts. In the case of an insurance contract with scaled payments, we construct estimators for the first moment of all future payments and present a proof of concept with a numerical example using simulated data. This is the basis of any type of risk assessment, where the state process of the insured follows the non-restrictive non-Markov model. Furthermore, estimating bivariate conditional transition rates and transition probabilities allows the estimation of second moments of future liabilities with a canonical one-dimensional cash flow in the non-Markov model. Altogether, this can be used to evaluate the widespread use of Markov assumptions in the reserving and premium calculation for insurance portfolios.
@@
Acknowledgements
The author would like to express gratitude to Marcus Christiansen and Christian Furrer for their invaluable discussions regarding the direction and theoretical background of this paper.
@@ Appendix A Appendix
For a monotonically increasing real-matrix valued function $\Lambda$ and a function $\phi(\boldsymbol{t})$, the equation
$Y(\boldsymbol{t})=\phi(\boldsymbol{t})+\int\displaylimits_{(s,\boldsymbol{t}]}Y(\boldsymbol{u}^{-})\Lambda(\mathrm{d}\boldsymbol{u}),$ (A.1)
@@
$\begin{split}\tilde{P}(\boldsymbol{t})&=\phi(\boldsymbol{t})+\int\displaylimits_{(s,\boldsymbol{t}]}\phi(\boldsymbol{u}^{-})\Lambda(\mathrm{d}\boldsymbol{u})\mathcal{P}((\boldsymbol{u},\boldsymbol{t}],\Lambda),\end{split}$ (A.2)
@@
For monotonically increasing functions $A,B:\mathbb{R}^{2}\rightarrow\mathbb{R}^{n\times n}$ we have the following equation
@@
Additionally, we need a two-dimensional generalisation of the well known Campbell theorem, see for example Milbrodt & Helbig [22].
@@
For the estimation of integrals, we use the so called one and two-dimensional variation of a process. We do a short introduction here.